OpenStudy (anonymous):

how to convert 33.2 kg/l to kg/ml? show work? by cancelling

OpenStudy (phi):

you have \( \frac{kg}{L} \) and you want to change the L to mL if you think "I want to cancel the L" and replace it with mL then multiply by \( \frac{L}{mL} \) pretend L is a number , then when we do \[ \frac{kg}{L}\cdot \frac{L}{mL} = \frac{kg}{\cancel{L}}\cdot \frac{\cancel{L}}{mL} = \frac{kg}{mL}\]

OpenStudy (phi):

to use that idea you want the actual number of L per ml there are 1000 ml per L \[ 1000 \frac{mL}{L}\] but we want the "flipped version" \[ \frac{1}{1000} \frac{L}{mL} \] multiply that times your number \[ 33.2 \frac{kg}{L} \cdot \frac{1}{1000} \frac{L}{mL} \]